OpenStudy (anonymous):

solve the equation for exact solutions over the interval [0,2pi). (cscx+2)(cscx-√2)=0

OpenStudy (anonymous):

Id love to help but i do not know this stuff. Sorry!

OpenStudy (campbell_st):

well the equation has been factored for you.. and csc(x) = 1/sin(x) so making the substitution you need to solve 1/sin(x) + 2 = 0 so 1/sin(x) = -2 take the reciprocal sin(x) = -1/2 similar for the other factor you will have 2 exact values... and you can find them in the given domain. hope it helps

OpenStudy (anonymous):

it does thank you... the square root over 2 was just stumping me

OpenStudy (anonymous):

pi/4 and pi/6 ??